# Dynamics and modulation of cosmic ray modified magnetosonic waves in a galactic gaseous rotating plasma

### v2

@@ V.1 Rogue Wave solutions
The MI of the magnetosonic wavepackets leads to the generation of freak waves. Generations of such waves, in the region $MN>0$, are noticed to appear instantaneously and depart without any track down. Subsequently, the NLS Eq. (46) admits rational solutions localized in both space and time variables kharif2008rogue ; moslem2011dust as follows,
$B_{1}(\xi,\tau)=\sqrt{M/N}\left[\frac{4(1+2iM\tau)}{1+4\xi^{2}+4M^{2}\tau^{2}}-1\right]\exp(iM\tau).$ (53)
The first-order rational rogue solution (53) reveals that a significant amount of magnetosonic wave energy is concentrated in a relatively small area in space. Typically, rogue waves are the envelopes of carrier waves that have wavelengths smaller than those of the central region of the envelope. These waves have the unusual feature of not being periodic in both space and time. Combinations of multiple first-order rogue waves by nonlinear superpositions lead to higher-order rogue wave solutions having more complicated nonlinear forms and higher amplitude. The amplitude of such waves is generally four to five times the amplitude of the background wave. The existence of second-order rogue waves was experimentally observed by Chabchoub et al. chabchoub2012super in surface water gravity waves. The exact second-order rogue wave solution of the NSL equation (46) is given by akhmediev2009rogue ,
$B_{2}(\xi,\tau)=\sqrt{M/N}\left(1+\frac{M_{2}+iN_{2}}{O_{2}}\right)\exp(iM\tau),$ (54)
@@
$M_{2}=\frac{3}{8}-\frac{1}{2}\xi^{4}-\frac{3}{2}\xi^{2}-6(P\xi\tau)^{2}-10(P\tau)^{4}-9(P\tau)^{2},$
$N_{2}=-P\tau\left[-\frac{15}{4}+\xi^{4}-3\xi^{2}+4(P\xi\tau)^{2}+4(P\tau)^{4}+2(P\tau)^{2}\right],$
$O_{2}=\frac{3}{32}+\frac{1}{12}\xi^{6}+\frac{1}{8}\xi^{4}+\frac{1}{2}\xi^{4}(P\tau)^{2}+\frac{9}{16}\xi^{2}+\xi^{2}(P\tau)^{4}$
$-\frac{3}{2}(P\xi\tau)^{2}+\frac{2}{3}(P\tau)^{6}+\frac{9}{2}(P\tau)^{4}+\frac{33}{8}(P\tau)^{2}.$
@@
Figure 10: A comparison between first- and second-order rogue wave pulses $|B_{1}|,|B_{2}|$ is displayed against $\xi$. Here, $(C_{g},C_{c},\Omega_{0},\theta)=(0.5,0.2,0.9,\pi/3)$ and $(C_{g},C_{c},\Omega_{0},\theta)=(0.1,0.7,0.9,\pi/3)$ in subplots (a) and (b), respectively. All other parameters remain unchanged as in Fig. 2.
@@
To gain some insight, we analyse the effects of the parameters $C_{g}$ and $C_{c}$ related to thermal and cosmic ray pressures, respectively, as well as the impact of the Coriolis force due to the rotation of the cosmic fluids with the angular frequency ($\Omega_{0}$) and the angle of rotation ($\theta$), on both the first- and second-order rogue wave solutions (53) and (54). Initially, the typical first-order rogue wave profiles $|B_{1}|$ are plotted against $\xi$ and $\tau$ for two different values of $\Omega_{0}$ in Fig. 7, keeping all other parameters fixed as mentioned in the figure caption. It has been observed that the amplitude and width of the rogue wave profile are reduced due to the enhancement of rotational frequency. The observation is quite natural, as the pulse can not gain energy from the background wave. To provide more specific details on the effects of the essential parameters on the evaluation of rogue waves, we plot the first-order rogue wave profiles $|B_{1}|$ against $\xi$ at an instant $\tau=0$ for different values of $\Omega_{0}$, $\theta$ (subplots (a) in Fig. 8) and $C_{g}$, $C_{c}$ (subplots (b) of Fig. 8). The figures illustrate that as the values of the parameters $\Omega_{0}$ and $\theta$ increase both the amplitude and width decrease, i.e., an increase in $\Omega_{0}$ and/or $\theta$ leads to the contraction of the rogue wave pulse. On the other hand, we notice that the inclusion of cosmic ray pressure with the thermal pressure significantly enlarges the rogue wave pulses. Subplot (b) in Fig. 8 depicts that both the amplitude and width increase with the increasing value of $C_{g}$ and $C_{c}$ related to the thermal pressure and cosmic ray pressure, respectively. These observations in Fig. 8) are found to be similar for the second-order solution $|B_{2}|$ versus $\xi$ in $\tau=0$ as displayed in Fig. 9. However, the second-order pulses’ amplitude is higher than that of first-order rogue wave pulses. The reduction (enhancement) of the amplitude and width of the rogue wave profiles, for both the first- and second-order solutions, signifies that the nonlinearity of the system diminishes (increases) due to the increase in the values of $\Omega_{0}$ and/or $\theta$ ($C_{g}$ and/or $C_{c}$ ). Hence, the Coriolis force due to the rotation of the cosmic fluids, and the modified pressure due to the inclusion of cosmic ray pressure with thermal pressure exhibit stabilizing behaviour on first- and second-order rogue wave solutions.
@@
VI CONCLUSIONS
@@
To conclude, we have investigated the evaluation of linear and nonlinear magnetosonic waves in the interstellar medium of spiral galaxies, which are the combination of thermal and cosmic fluid, taking into account the magneto-rotational effects of the fluid. The typical consistent parameters adopted in our numerical assessment are relevant in such fluid mediums of spiral galaxies gliddon1966gravitational ; turi2022magnetohydrodynamic . The cosmic fluid is found to have a reasonable contribution in modifying the thermal pressure in terms of the system’s total pressure. Consequently, the dispersion properties, nonlinear wave evaluation, and modulation instabilities are modified. The effects of the parameters $C_{g}$ and $C_{c}$, related to the thermal and cosmic ray pressure, respectively, have been demonstrated. The Coriolis force due to the rotation of the plasma is found to have significant refinement in the wave propagation as the rotation frequency ($\Omega_{0}$) and rotation angle ($\theta$) vary. Further, cosmic ray diffusivity ($\kappa$) and magnetic resistivity ($\eta$) play a crucial role in evaluating shock waves. The significant results obtained are summarized as follows:
@@
  • The linear analysis reveals that the dispersion relation (III) remains the same as obtained in Turi and Misraturi2022magnetohydrodynamic in the absence of both cosmic ray diffusivity ($\kappa$) and magnetic resistivity ($\eta$). The waves become unstable due to the increase in the values of the parameters $\eta$ and $\kappa$. However, interestingly, the damping caused by cosmic ray diffusivity and magnetic resistivity becomes negligible as one increases the rotating frequency. Wave damping is reduced by increasing the effects of the rotational effect in the system. The damping is also modified due to the presence of cosmic fluid in the system by accelerating the damping rate $|\gamma|$.
@@
  • To perform nonlinear analysis, we first derived a KdVB equation by employing the reductive perturbation technique. In solving the KdVB equation, several nonlinear wave shapes have been evaluated, analytically, and numerically. The results indicate that the parameters of cosmic ray diffusivity ($\kappa_{0}$) and magnetic resistivity ($\eta_{0}$) are responsible for the formation of shock structures in the current model. When the effects of $\kappa_{0}$ and $\eta_{0}$ grow considerably, the dispersion effect becomes negligible compared to dissipation, and the monotonic shocks become stationary. Additionally, it has been noticed that the numerical solutions lead to oscillatory shock wave profiles, and growing values of $\kappa_{0}$, $\eta_{0}$ and rotation coefficients produce damped oscillatory waves with reduced amplitude. The KdVB equation becomes the KdV equation without the presence of $\eta_{0}$ and $\kappa_{0}$, consequently, the solitary wave evaluation has been reported.
@@
  • To characterize the weakly nonlinear development of the envelope of a modulated wave packet in the low-frequency limit, we derive a nonlinear Schrödinger (NLS) equation from Eq. (21). It has been noticed that incorporating cosmic ray pressure in the present model reduces MI growth. As the modulated wave number $K$ increases, the growth rate rises until it hits a critical growth rate $\Gamma_{c}$, which varies depending on the plasma properties. It sharply declines for even higher values of $K$ after hitting the critical value growth rate. Furthermore, the present model supports the propagation of magnetosonic rogue waves. It is observed that the Coriolis force reduces the system’s nonlinearity, resulting in a shorter rogue wave amplitude, however, the inclusion of cosmic ray pressure with the thermal pressure significantly enlarges the rogue wave pulses.
@@
Our investigation is based on typical plasma parameters: $B_{0}\sim 1$ nT, $\rho_{0}\sim 10^{-21}$ Kg/m3, $P_{g0}\sim 10^{-13}$ N/m2, $P_{c0}\sim 10^{-13}$/m2, $\eta\sim 10^{-1}V_{A}^{2}/\omega_{ci}$ and $\kappa\sim 10^{-2}V_{A}^{2}/\omega_{ci}$ which are relevant to the spiral galaxies. gliddon1966gravitational ; turi2022magnetohydrodynamic Thus, we believe the present study could be useful for understanding the magnetosonic wave propagation in the interstellar medium of spiral galaxies, where cosmic rays play an important role.
@@ ACKNOWLEDGEMENTS
One of us, J. Turi is grateful to the Council of Scientific and Industrial Research (CSIR) for a Senior Research Fellowship (SRF) with Ref. No. 09/202(0115)/2020-EMR-I.

### v3

@@ V.1 Rogue Wave Solutions
The MI of the magnetosonic wavepackets leads to the generation of freak waves. Generations of such waves in the region $MN>0$, are noticed to appear instantaneously and depart without any track down. Subsequently, the NLS Eq. (46) admits rational solutions localized in both space and time variables kharif2008rogue ; moslem2011dust as follows,
$B_{1}(X,T)=\sqrt{M/N}\left[\frac{4(1+2iMT)}{1+4X^{2}+4M^{2}T^{2}}-1\right]\exp(iMT).$ (53)
The first-order rational rogue solution (53) reveals that a significant amount of magnetosonic wave energy is concentrated in a relatively small area in space. Typically, rogue waves are the envelopes of carrier waves having wavelengths smaller than those of the central region of the envelope. These waves have the unusual feature of not being periodic in both space and time. Combinations of multiple first-order rogue waves by nonlinear superpositions lead to higher-order rogue wave solutions having more complicated nonlinear forms and higher amplitude. The amplitude of such waves is generally four to five times the amplitude of the background wave. The existence of second-order rogue waves was experimentally observed by Chabchoub et al. chabchoub2012super in surface water gravity waves. The exact second-order rogue wave solution of the NLS equation (46) is given by akhmediev2009rogue ,
$B_{2}(X,T)=\sqrt{M/N}\left(1+\frac{M_{2}+iN_{2}}{O_{2}}\right)\exp(iMT),$ (54)
@@
$M_{2}=\frac{3}{8}-\frac{1}{2}X^{4}-\frac{3}{2}X^{2}-6(PXT)^{2}-10(PT)^{4}-9(PT)^{2},$
$N_{2}=-PT\biggr{[}-\frac{15}{4}+X^{4}-3X^{2}+4(PXT)^{2}+4(PT)^{4}$
$+2(PT)^{2}\biggr{]},$
$O_{2}=\frac{3}{32}+\frac{1}{12}X^{6}+\frac{1}{8}X^{4}+\frac{1}{2}X^{4}(PT)^{2}+\frac{9}{16}X^{2}+X^{2}(PT)^{4}$
$-\frac{3}{2}(PXT)^{2}+\frac{2}{3}(PT)^{6}+\frac{9}{2}(PT)^{4}+\frac{33}{8}(PT)^{2}.$
@@
Figure 10: A comparison between first- and second-order rogue wave pulses $|B_{1}|,|B_{2}|$ is displayed [see Eqs. (53) and (54)] against $X$ for different values of the rotational frequency $\Omega_{0}$ and angle of rotation $\theta$. The second-order correction of the rogue wave solution enhances the amplitude and width of the wave pulses. Compared to first-order rogue waves, second-order rogue waves are spikier because they absorb more energy from the surrounding waves. Here, $(C_{g},C_{c},\Omega_{0},\theta)=(0.5,0.2,0.9,\pi/3)$ and $(C_{g},C_{c},\Omega_{0},\theta)=(0.1,0.7,0.9,\pi/3)$ in subplots (a) and (b), respectively. All other parameters remain unchanged as in Fig. 2.
@@
To gain some insight, we analyse the effects of the parameters $C_{g}$ and $C_{c}$ related to thermal and cosmic ray pressures, respectively, as well as the impact of the Coriolis force due to the rotation of the cosmic fluids with the angular frequency ($\Omega_{0}$) and the angle of rotation ($\theta$), on both the first- and second-order rogue wave solutions (53) and (54). Initially, the typical first-order rogue wave profiles $|B_{1}|$ are plotted against $X$ and $T$ for two different values of $\Omega_{0}$ in Fig. 7, keeping all other parameters fixed as mentioned in the figure caption. It has been observed that the amplitude and width of the rogue wave profile are reduced due to the enhancement of rotational frequency. The observation is quite natural, as the pulse cannot gain energy from the background wave. To provide more specific details on the effects of the essential parameters on the evaluation of rogue waves, we plot the first-order rogue wave profiles $|B_{1}|$ against $X$ at an instant $T=0$ for different values of $\Omega_{0}$, $\theta$ (subplots (a) in Fig. 8) and $C_{g}$, $C_{c}$ (subplots (b) of Fig. 8). The figures illustrate that as the values of the parameters $\Omega_{0}$ and $\theta$ increase both the amplitude and width decrease, i.e., an increase in $\Omega_{0}$ and/or $\theta$ leads to the contraction of the rogue wave pulse. On the other hand, we notice that the inclusion of cosmic ray pressure with the thermal pressure significantly enlarges the rogue wave pulses. Subplot (b) in Fig. 8 depicts that both the amplitude and width increase with the increasing value of $C_{g}$ and $C_{c}$ related to the thermal pressure and cosmic ray pressure, respectively. The results obtained indicate that, in the ISM of spiral galaxies, the temporal evaluation of rogue waves gains a significant amount of energy due to the modification of thermal pressure by cosmic ray pressure. Increasing rotational frequency leads to dissipation of energy and decreases the rogue wave amplitude. The influence of cosmic rays amplified the rogue wave, which indicates the creation of regions of an enhanced magnetic field. These observations in Fig. 8 are found to be similar for the second-order solution $|B_{2}|$ versus $X$ at $T=0$ as displayed in Fig. 9. However, the second-order pulses’ amplitude is higher than that of first-order rogue wave pulses. The reduction (enhancement) of the amplitude and width of the rogue wave profiles, for both the first- and second-order solutions, signifies that the nonlinearity in the present model diminishes (increases) due to the increase in the values of $\Omega_{0}$ and/or $\theta$ ($C_{g}$ and/or $C_{c}$ ). Hence, the Coriolis force due to the rotation of the cosmic fluids, and the modified pressure due to the inclusion of cosmic ray pressure with thermal pressure exhibit stabilizing behaviour on first- and second-order rogue wave solutions.
@@
VI SENSITIVITY ANALYSIS
In the present study, we have investigated the excitations of nonlinear magnetosonic solitons, shocks, and rouge waves in the ISM of spiral galaxies. The analytical and numerical studies of these nonlinear structures, presented in Secs. IV and V, imply that the evaluation and characteristics of such waves largely depend on the parameters described in this model. We have observed from our obtained numerical results that a small change in some parameters may lead to a significant influence on the wave characteristics. In order to study the relative importance of different plasma parameters and identify the parameters that influence the magnetic field strength $B$ most, we perform a normalized forward sensitivity analysis dwivedi2022modeling ; mandal2023dynamic . To investigate sensitivity analysis, we compute the normalized forward sensitivity indices $S_{\mu}^{B}$ of $B$ for a parameter $\mu$ (where $\mu$ is either of the parameters $C_{g},C_{c},\Omega_{0},\theta,\kappa_{0},\eta_{0}$), as given below,
$S_{\mu}^{B}=\frac{\mu}{B}\times\frac{\partial B}{\partial\mu}.$ (55)
We calculate $S_{\mu}^{B}$ for each plasma parameter at the typical normalized parametric value $(C_{g},C_{c},\Omega_{0},\theta,\kappa_{0},\eta_{0})$ $=(0.7,0.6,0.8,\pi/7,0.9,0.02)$. The calculated values of the magnetic field strength sensitivity index corresponding to the soliton solution (26), the shock solution (38), and the first-order rogue wave solution (53) at $(\zeta,v_{0})=(0.8,0.1)$, $\chi=50$, and $(X,T,k)=(0.4,0.2,0.4)$, respectively, are presented in Table II. When the sensitivity index $S_{\mu}^{B}$ is positive, it means that the magnetic field strength $B$ is increasing as a function of the related parameter $\mu$, when it is negative, it means that $B$ is decreasing. To illustrate, we obtain the sensitivity index $S_{C_{c}}^{B}=2.2748$ of $B$ for the parameter $C_{c}$ corresponding to the magnetosonic shocks, which represents an increase of $1\%$ in $C_{c}$ will increase the magnetic field $B$ by $2.2748\%$. However, $S_{C_{g}}^{B}=-2.0590$ indicates that $B$ decreases by $2.0590\%$ as $C_{g}$ increases by $1\%$. The values obtained in Table II reveal that, for the magnetosonic shock structure, $C_{c}$ appears to be the most sensitive parameter with a positive impact on the magnetic field strength $B$. The parameters $C_{g}$ and $\theta$ have a negative sensitive impact on $B$ corresponding to the shock solution, and all other parameters have a positive impact. The parameters $C_{g}$, $C_{c}$, $\Omega_{0}$, and $\theta$ have negative influences on the magnetic field strength of the magnetosonic soliton with $\theta$ as the most sensitive parameter. However, we find that these parameters have positive influences on the magnetic field strength of the magnetosonic rogue wave, of which $\Omega_{0}$ is the most sensitive parameter.
Sensitivity index ($S_{\mu}^{B}$) corresponding to
Physical Parameters ($\mu$) Soliton solution Shocks solution Rogue solution
Thermal speed ($C_{g}$) $-0.2671$ $-2.0590$ $0.6774$
Cosmic ray speed ($C_{c}$) $-0.1966$ $2.2748$ $0.4977$
Rotational frequency ($\Omega_{0}$) $-0.3912$ $2.0109$ $2.1591$
Rotational angle ($\theta$) $-1.9347$ $-1.4392$ $0.4576$
Magnetic resistivity ($\eta$) - $0.1169$
Cosmic ray diffusivity ($\kappa$) - $1.8379$
Table 2: Calculated sensitivity indices ($S_{\mu}^{B}$) of magnetic field strength $B$ with respect to the physical parameters $\mu=C_{g},C_{c},\Omega_{0},\theta,\kappa_{0},\eta_{0}$ corresponding to the magnetosonic soliton, shock, and rogue wave solutions are displayed at the typical normalized parametric value $(C_{g},C_{c},\Omega_{0},\theta,\kappa_{0},\eta_{0})$ $=(0.7,0.6,0.8,\pi/7,0.9,0.02)$.
VII CONCLUSIONS
@@
To conclude, we have investigated the evaluation of linear and nonlinear magnetosonic waves in the interstellar medium of spiral galaxies, which are the combination of thermal and cosmic fluid, taking into account the magneto-rotational effects of the fluid. The typical consistent parameters adopted in our numerical assessment are relevant in such fluid mediums of spiral galaxies gliddon1966gravitational ; turi2022magnetohydrodynamic . The cosmic fluid is found to have a reasonable contribution in modifying the thermal pressure in terms of the total pressure. Consequently, the dispersion properties, nonlinear wave evaluation, and modulation instabilities are modified. The effects of the parameters $C_{g}$ and $C_{c}$, related to the thermal and cosmic ray pressure, respectively, have been demonstrated. The Coriolis force due to the rotation of the plasma is found to have significant refinement in the wave propon frequency ($\Omega_{0}$) and rotation angle ($\theta$) vary. Further, cosmic ray diffusivity ($\kappa$) and magnetic resistivity ($\eta$) play a crucial role in evaluating shock waves. The significant results obtained are summarized as follows:
@@
  • The linear analysis reveals that the dispersion relation (III) remains the same as obtained in Turi and Misraturi2022magnetohydrodynamic in the absence of both cosmic ray diffusivity ($\kappa$) and magnetic resistivity ($\eta$). The waves become unstable due to the increase in the values of the parameters $\eta$ and $\kappa$. However, interestingly, the damping caused by cosmic ray diffusivity and magnetic resistivity becomes negligible as one increases the rotating frequency. Wave damping is reduced by increasing the effects of the rotational effect in the plasma model. The damping is also modified due to the presence of cosmic fluid in the ISM of spiral galaxies by accelerating the damping rate $|\gamma|$.
@@
  • To perform nonlinear analysis, we first derived a KdVB equation by employing the reductive perturbation technique. In solving the KdVB equation, several nonlinear wave shapes have been evaluated analytically, and numerically. The results indicate that the parameters of cosmic ray diffusivity ($\kappa_{0}$) and magnetic resistivity ($\eta_{0}$) are responsible for the formation of shock structures in the current model. When the effects of $\kappa_{0}$ and $\eta_{0}$ grow considerably, the dispersion effect becomes negligible compared to dissipation, and the monotonic shocks become stationary. Additionally, it has been noticed that the numerical solutions lead to oscillatory shock wave profiles, and growing values of $\kappa_{0}$, $\eta_{0}$ and rotation coefficients produce damped oscillatory waves with reduced amplitude. The KdVB equation becomes the KdV equation without the presence of $\eta_{0}$ and $\kappa_{0}$, consequently, the solitary wave evaluation has been reported.
@@
  • To characterize the weakly nonlinear development of the envelope of a modulated wave packet in the low-frequency limit, we derive a nonlinear Schrödinger (NLS) equation from Eq. (21). It has been noticed that incorporating cosmic ray pressure in the present model reduces MI growth. As the modulated wave number $K$ increases, the growth rate rises until it hits a critical growth rate $\Gamma_{c}$, which varies depending on the plasma properties. It sharply declines for even higher values of $K$ after hitting the critical value growth rate. Furthermore, the present model supports the propagation of magnetosonic rogue waves. It is observed that the Coriolis force reduces the nonlinearity of the model, resulting in a shorter rogue wave amplitude, however, the inclusion of cosmic ray pressure with the thermal pressure significantly enlarges the rogue wave pulses.
@@
Our investigation is based on typical plasma parameters: $B_{0}\sim 1$ nT, $\rho_{0}\sim 10^{-21}$ Kg/m3, $P_{g0}\sim 10^{-13}$ N/m2, $P_{c0}\sim 10^{-13}$/m2, $\eta\sim 10^{-2}V_{A}^{2}/\omega_{ci}$ and $\kappa\sim 10^{-1}V_{A}^{2}/\omega_{ci}$ which are relevant to the spiral galaxies gliddon1966gravitational ; turi2022magnetohydrodynamic (See. Table I for more detail). The obtained results show that cosmic rays modify the existing pressure law and enhance the galactic magnetic field strength due to the interaction of cosmic rays with ionizied gas in the ISM clouds of spiral galaxies. The interplay between cosmic ray diffusivity and magnetic resistivity leads to the evaluation of shock waves in ISM of spiral galaxies. The higher values of these two parameters amplify the shock waves. Whereas, the lower values of these parameters lead to the formation of sloitons and rogue waves. The parameter related to the cosmic ray pressure in ISM clouds significantly modifies these wave structures. Thus, we believe the present study could be useful for understanding the magnetosonic wave propagation in the interstellar medium of spiral galaxies, where cosmic rays play an important role.
@@ ACKNOWLEDGEMENTS
One of us, J. Turi is grateful to the Council of Scientific and Industrial Research (CSIR) for a Senior Research Fellowship (SRF) with Ref. No. 09/202(0115)/2020-EMR-I. The authors thank Ms. Snehalata Nasipuri and Mr. Gourav Mandal for the valuable discussion to improve the work. The authors also thank the anonymous referee for his crucial and constructive suggestions to build up the work in its present form.